# GP-MoLFormer: A Foundation Model For Molecular Generation

### v1

@@ Abstract
We compare the performance of GP-MoLFormer in terms of generating DRD2 active molecules that will pass the DRD2 binding classifier ($p>0.5$). For baselines of comparison, we consider our own random generations from GP-MoLFormer-Raw, as well as an earlier scaffold-conditioned generation model [26] that was specifically trained for scaffold decoration tasks and was then used to decorate the same scaffolds under investigation here with fragments from ChEMBL. In contrast to this baseline model, GP-MoLFormer-Raw has not seen scaffold-constrained generation during pre-training, nor is it specifically fine-tuned for this purpose. Table 6 shows that GP-MoLFormer-Raw generates more DRD2 active hits compared to a random baseline of de novo generation, as well as a generative model trained on this specific task.
@@
de novo GP-MoLFormer-Raw 0.83
scaffold-conditioned GP-MoLFormer-Raw 4.58
Table 6: Scaffold-constrained generation. Baseline performance is taken from Arús-Pous et al. (2020)[26].
@@
We exploit prompt-tuning to introduce a novel means for enabling GP-MoLFormer to tackle property-specific molecular optimization tasks, where the goal is to generate molecules with a specific property value above a pre-defined threshold. Below, we describe the pair-tuning framework and then show that pair-tuning performs well on a set of three tasks.
@@
1:Pretrained GP-MoLFormer; Number of fine-tuning epochs, $m$; number of enhancement tokens, $n$; Pair-Tuning dataset $D=\{(a,b)|b>a\}$ for $a,b\in\Omega$ where $>$ is an order relation defined for a certain property $T$
2:Append $n+1$ new tokens $\phi_{T}=\langle enh_{1}\rangle\langle enh_{2}\rangle\dots\langle enh_{n}\rangle,\langle sep\rangle$ to GP-MoLFormer-Raw vocabulary
@@
Pairtuning tasks
We evaluate pairtuning using GP-MoLFormer-Raw on three property optimization benchmarks, namely drug-likeness (QED) maximization, penalized logP maximization, and activity maximization for DRD2. The first two properties, QED and penalized logP, are important considerations in drug discovery, and these task shows the ability of a model to optimize salient aspects of a molecule, even if maximization of these properties by themselves is of low utility[30]. The goal of the third task is to increase the binding affinity of a compound to a ligand, in this case, the dopamine $\mathrm{D}_{2}$ receptor.
@@
Table 7: Performance on unconstrained penalized logP and QED optimization. Baseline performances are taken from Zhou et al. (2019)[30] and Eckmann et al. (2022)[20] and are reported on 100k generations as per LIMO[20]. For GP-MoLFormer, we set $k$ the number of targeted generation attempts per molecule to $125$: given a test set of size 800 this results in 100k total generations.
@@
Table 7 shows results of pair-tuning on GP-MoLFormer-Raw, as well as of the baselines, in terms of the generated molecules with high penalized logP. Penalized logP is calculated as $\log\mathrm{P}-\mathrm{SA}-{\max(\max_{\mathrm{rings}}(\mathrm{size})-6,0)}$, i.e., logP penalized by SA and maximum ring size, if larger than 6. We report pair-tuning performances as a function of two different $k$, where $k$ is the number of targeted generation attempts per molecule. For $k=125$, using a test set containing 800 molecules gives a total number of generated molecules of 100k, which is the same used for the baselines. The baselines under consideration are JT-VAE [19], GCPN [31], MolDQN [30], MARS [32], GraphDF [33], and LIMO [20]. Penalized logP can be artificially inflated simply by generating molecules with increased length, specifically by adding alkyl carbons [32, 20]. Many works, e.g., GCPN, MolDQN, and LIMO, avoid this by reporting top property scores given length constraints, e.g., limiting the length up to the maximum molecule size of the ZINC250k dataset. MARS, on the other hand, does not consider such a length constraint. We also report the top 3 scores for pair-tuning with a length constraint (length < 38), added post generation, in Table 7 as the value within parentheses.
Compared to the strongest length-constrained baselines, pair-tuning generates molecules with comparably high values. When the length constraint is not considered, pair-tuning still generates molecules with higher but reasonable penalized logP values. Note that pair-tuning does not require feedback or evaluation on generations from an additional reward model or a property predictor, nor is the generative model updated during the tuning. We also report top 3 scores for 1M generations ($k=1000$), which requires less than an hour to generate. Although all the baselines produce molecules with 100% validity due to their methods utilizing SELFIES or graphs, our method’s validity is still very high (around 95%) and overall this is negligible compared to the ease of generating additional molecules. Altogether, Table 7 shows that the proposed method can generate molecules with even higher penalized logP values, both with and without a length constraint.
@@
As with penalized logP, we show results for QED optimization in Table 7 compared with the same baselines. Again, pair-tuning performances are reported for two different values of $k$, showing comparable performances with respect to baselines.
@@
Table 8: Performance on the unconstrained DRD2 activity optimization task with respect to the initial value. Baseline performances are reported from Erikawa et al. (2023)[34].
@@
DRD2 activity optimization results are reported in Table 8. Activity scores are calculated using the trained predictor from Olivecrona et al. (2017)[35]. Average activity scores of the input seed molecules are also shown. Different baseline performances are reported using different test seed molecules, and we are interested in comparing the activity improvement of an experiment with the set of test seed molecules. For pair-tuning, performance reported considers $k=20$ generations per seed molecule and we use the top 1 for each, similar to the baselines, Mol-CycleGAN [36] — a graph-based generation method that uses a CycleGAN optimization scheme — and Gargoyles [34], which uses Monte Carlo Tree Search in fragment space to optimize molecular graphs based on an evaluation function. Results show that pair-tuning generates molecules with the highest activity improvement with respect to the seed molecules.
@@ Conclusion
In this work, we investigate the effect of training data bias and diversity on the downstream performances of a generative chemical language model named GP-MoLFormer, built on top of a recent transformer architecture for chemical language modeling. We show how bias in training data can induce memorization and thus impact the novelty of generation. To our knowledge, this is the first report demonstrating training data memorization and its impact on performance of a generative chemical language model pre-trained on billion-scale data. We further analyze how the commonly used metrics for comparing models’ generations with a held-out test set are affected by the diversity of the training distribution; hence, the sizes of the test set and the generation set should be carefully considered before making a conclusion based on those metrics. Finally, we show the generality of the proposed GP-MoLFormer architecture on two distinct targeted design tasks, i.e., scaffold-constrained molecular decoration and unconstrained property-guided molecular optimization, where the model performs better or on par when compared to baselines.
@@ Methods
The GP-MoLFormer decoder uses the transformer block used in MoLFormer. To avoid the quadratic complexity associated with regular attention computations in vanilla transformers [37], MoLFormer utilized a base 12-layer transformer architecture with linear attention [12], wherein each layer has $12$ attention heads and a hidden state size of $768$. A Generalized Feature map [12] for the linear attention was chosen.
@@
where $Q,K,V$ are the query, key, and value respectively, and $\varphi$ a random feature map. The GP-MoLFormer is trained on the next token prediction task using a cross-entropy objective: $L_{LM}(w1,...,wn)=\sum_{i}{\log{P}(w_{i}|w_{j<i})}$.
@@
We used two datasets for pre-training by combining SMILES from the PubChem [38] and ZINC [9] databases with varying proportions from each. The Raw dataset contains a total of 1.1B SMILES strings;  [^1] M of them are from the PubChem dataset, whereas the larger 1B portion comes from the ZINC database. The Clean dataset is a de-duplicated version of the Raw dataset, which comprises 650M SMILES. We utilized the tokenizer from Schwaller et al. (2019)[39] to construct a vocabulary. All SMILES sequences from both PubChem and ZINC are converted to a canonical format with no isomeric information using RDKit [40] followed by de-duplication (Clean only) and tokenization. All unique tokens extracted from the resulting output give us a vocabulary of $2357$ tokens plus $5$ special tokens, resulting in a total of $2362$ vocabulary tokens which are used for all pre-trained models considered in this paper, irrespective of pre-training dataset size. The post tokenization sequence length of the molecules range from $1$ to just over $2000$ tokens. We decide to restrict the sequence length range from $1$ token to $202$ tokens, special tokens inclusive, to reduce computation time. Since over $99.4$ percent of all molecules from our dataset contain less than $202$ tokens, we hypothesize that the removal of molecules with more than $202$ tokens would be of minimal negative impact on pre-training.
Large Scale Training and Parallelization
For pre-training, we use the causal language model objective defined in Devlin et al. (2019)[41]. The training was performed for 4 epochs through the entire combined PubChem+ZINC dataset with a fixed learning rate of $1.6\times 10^{-4}$ and a batch size of $1600$ molecules per GPU on a total of $16$ A100 80 GB GPUs over $2$ servers connected via EDR Infiniband fabric. It should be noted that as the number of GPUs utilized increased, we found an increase in learning rate was necessary up to a factor of $8$.
@@
For evaluation, we do a forward pass using the following sequence: the first 20 prompt embeddings + the <bos> embedding + the input molecule embeddings + the <unk> embedding. After that, we sample from the token distribution generated by GP-MoLFormer until <eos> is encountered. For all pair-tuning experiments, batch size was set to 35, the learning rate was fixed at $3\times 10^{-2}$, and the number of epochs run was 1000. Each epoch took 6 minutes to complete on a single GPU. During pairtuning generation, GP-MoLFormer takes $0.00375$ seconds for a single forward pass.

### v2

@@ Abstract
We compare the performance of GP-MoLFormer in terms of generating DRD2 active molecules that will pass the DRD2 binding classifier ($p>0.5$). For baselines of comparison, we consider our own random generations from GP-MoLFormer, as well as an earlier scaffold-conditioned generation model [26] that was specifically trained for scaffold decoration tasks and was then used to decorate the same scaffolds under investigation here with fragments from ChEMBL. In contrast to this baseline model, GP-MoLFormer has not seen scaffold-constrained generation task during pre-training, nor is it specifically fine-tuned for this purpose. Table 3 shows that GP-MoLFormer generates more DRD2 active hits compared to a random baseline of de novo generation, as well as a generative model trained on this specific task. Examples of scaffold-decorated molecules using GP-MoLFormer are shown in the Supplementary Figure S4.
@@
de novo GP-MoLFormer 0.83
scaffold-conditioned GP-MoLFormer 4.58
Table 3: Scaffold-constrained generation. Predicted active hits is the percentage of generated molecules that pass the DRD2 binding classifier. Baseline performance is taken from Arús-Pous et al. (2020)[26]. Bold value indicates the best performing model.
@@
We exploit prompt-tuning to introduce a novel means for enabling GP-MoLFormer to tackle property-specific molecular optimization tasks, where the goal is to generate molecules with a specific property value above a pre-defined threshold. Below, we describe the pair-tuning framework and then show that pair-tuning performs well on a set of three tasks. We evaluate pair-tuning using GP-MoLFormer on three property optimization benchmarks, namely drug-likeness (QED) maximization, penalized logP maximization, and activity maximization for DRD2. The first two properties, QED and penalized logP, are important considerations in drug discovery, and these task shows the ability of a model to optimize salient aspects of a molecule, even if maximization of these properties by themselves is of low utility[28]. The goal of the third task is to increase the binding affinity of a compound to a ligand, in this case, the dopamine $\mathrm{D}_{2}$ receptor.
@@
1:Pre-trained GP-MoLFormer; Number of fine-tuning epochs, $m$; number of enhancement tokens, $n$; Pair-Tuning dataset $D=\{(a,b)|b>a\}$ for $a,b\in\Omega$ where $>$ is an order relation defined for a certain property $T$
2:Append $n+1$ new tokens $\phi_{T}=\langle enh_{1}\rangle\langle enh_{2}\rangle\dots\langle enh_{n}\rangle,\langle sep\rangle$ to GP-MoLFormer vocabulary
@@
Table 4: Performance on unconstrained penalized logP and QED optimization. Pair-tuning is performed using frozen GP-MoLFormer. Baseline performances are taken from Zhou et al. (2019)[28] and Eckmann et al. (2022)[16] and are reported on 100k generations as per LIMO[16]. For GP-MoLFormer, we set $k$, the number of targeted generation attempts per molecule, to $125$ — given a test set of size 800 this results in 100k total generations. Values in parentheses are after post hoc length filtering. Bold values indicate the highest property values found (both length-constrained and unlimited).
@@
Table 4 shows results of pair-tuning on GP-MoLFormer, as well as of the baselines, in terms of the generated molecules with high penalized logP. Penalized logP is calculated as $\log\mathrm{P}-\mathrm{SA}-{\max(\max_{\mathrm{rings}}(\mathrm{size})-6,0)}$, i.e., logP penalized by SA and maximum ring size, if larger than 6. We report pair-tuning performances as a function of two different $k$, where $k$ is the number of targeted generation attempts per molecule. For $k=125$, using a test set containing 800 molecules gives a total number of generated molecules of 100k, which is the same used for the baselines. The baselines under consideration are JT-VAE [15], GCPN [31], MolDQN [28], MARS [32], GraphDF [33], and LIMO [16]. Penalized logP can be artificially inflated simply by generating molecules with increased length, specifically by adding alkyl carbons [32, 16]. Many works, e.g., GCPN, MolDQN, and LIMO, avoid this by reporting top property scores given length constraints, e.g., limiting the length up to the maximum molecule size of the ZINC250k dataset [34]. MARS, on the other hand, does not consider such a length constraint. We also report the top 3 scores for pair-tuning with a length constraint (length < 38), added post generation, in Table 4 as the value within parentheses. Compared to the strongest length-constrained baselines, pair-tuning generates molecules with comparably high values. When the length constraint is not considered, pair-tuning still generates molecules with higher but reasonable penalized logP values. Note that pair-tuning does not require feedback or evaluation on generations from an additional reward model or a property predictor, nor is the generative model updated during the tuning. We also report top 3 scores for 1M generations ($k=1000$), which requires less than an hour to generate. Although all the baselines produce molecules with 100% validity due to their methods utilizing SELFIES or graphs, our method’s validity is still very high (around 95%) and overall this is negligible compared to the ease of generating additional molecules. Altogether, Table 4 shows that the proposed method can generate molecules with even higher penalized logP values, both with and without a length constraint.
@@
As with penalized logP, we show results for QED optimization in Table 4 (also see Supplementary Figure S5 for generated molecules) compared with the same baselines. Again, pair-tuning performances are reported for two different values of $k$, showing comparable performances with respect to baselines. Supplementary Tables S2 and S3 further demonstrate that pair-tuning with GP-MoLFormer produces higher scoring molecules that also share high diversity as well as high closeness to training distribution, compared to baselines.
@@
Table 5: Performance on the unconstrained DRD2 activity optimization task with respect to the initial value. Baseline performances are reported from Erikawa et al. (2023)[35]. Bold value indicates the best performing model.
@@
DRD2 activity optimization results are reported in Table 5. Activity scores are calculated using the trained predictor from Olivecrona et al. (2017)[36]. Average activity scores of the input seed molecules are also shown. Different baseline performances are reported using different test seed molecules, and we are interested in comparing the activity improvement of an experiment with the set of test seed molecules. For pair-tuning, performance reported considers $k=20$ generations per seed molecule and we use the top 1 for each, similar to the baselines, Mol-CycleGAN [37] — a graph-based generation method that uses a CycleGAN optimization scheme — and Gargoyles [35], which uses Monte Carlo Tree Search in fragment space to optimize molecular graphs based on an evaluation function. Results show that pair-tuning generates molecules with the highest activity improvement with respect to the seed molecules, when compared to baselines. Examples of generated molecules using pair-tuning on GP-MoLFormer are provided in the Supplementary Figure S6.
Scaling results
Effect of training dataset and generation pool
The generative ability of a chemical language model can be affected by the scale of the training data. Further, bias in training data can also contribute to the quality of generation. To disentangle these factors, we report GP-MoLFormer’s performance on two independently varying dimensions: First, we vary the size and quality of pre-training data. For this dimension, we compare GP-MoLFormer trained on 1.1B SMILES extracted from ZINC and PubChem, and GP-MoLFormer-Uniq trained on a 650M de-duplicated subset of that 1.1B SMILES set. Secondly, we analyze the effect of varying the number of generated molecules from 30k to 10B molecules, to study the effect of inference time scaling.
To summarize, GP-MoLFormer is trained on a dataset of 650M–1.1B SMILES, which captures the relative abundance of molecules, as well as the presence of the same molecule in different context, as found in chemical databases and is evaluated on generations up to a scale of billions. This is in contrast to the existing molecular generation benchmarks that report performance metrics for a relatively small 10–30k generations, and to the current generative molecular models that are designed to target a specific distribution of molecules, e.g., synthetic molecules with biological activity or natural products, and are trained on 1–100M samples[14].
[…]
Additionally, when comparing the 10B molecules generated by the GP-MoLFormer and by the GP-MoLFormer-Uniq model, 67 to 74% of the novel molecules generated by a model are unique to that model (i.e., not in the other model’s generated set). This implies that the two models learned separate but overlapping manifolds. This aspect of different coverage of the molecular manifold with different model variants will be investigated further in future work.
This result confirms that (i) GP-MoLFormer trained on a billion of SMILES memorizes training samples, as seen from the high number of exact matches ($1-$novelty, which can be up to 60%) with training molecules; and (ii) training memorization becomes less when the training data is de-duplicated, enabling more novel generation. (iii) With scaling of inference compute, novelty in generations reduces, but remains significant, even when evaluated against $\approx$10B generations. In summary, in all cases studied here, GP-MoLFormer is capable of generating novel, diverse, and valid molecules.
Training Size = 650M Training Size = 1.1B
Generation Size Novel Unique Valid Novel Unique Valid
30k 0.390 0.997 1.000 0.323 0.997 0.997
100k 0.393 0.996 0.999 0.326 0.998 0.998
1M 0.395 0.996 0.999 0.323 0.996 0.997
10M 0.400 0.991 0.996 0.322 0.989 0.997
100M 0.385 0.947 0.996 0.327 0.989 0.997
1B 0.340 0.675 0.996 0.278 0.611 0.997
10B 0.214 0.270 0.996 0.167 0.223 0.997
Table 6: Novelty, validity, and uniqueness of different number of generations for GP-MoLFormer trained with 650M (Uniq) and 1.1B (Raw) training set.
Discussions on the effect of training data bias on generations
Within the 1.1B training set, a notable 45% of SMILES strings were found duplicates. Some of this is due to molecules that appear multiple times as different isomers, which are pre-processed into the same SMILES representation in our canonicalization pipeline. There is also some overlap between the two databases — ZINC and PubChem. This popularity of certain molecules may or may not be related to true factors, such as the molecule being useful in a multitude of applications because of its synthetic ease or lower cost, or a combination of all of those. It is reported that such skewed distribution can also originate from anthropogenic attention bias [11], where some molecules are studied extensively because scientists themselves or their peers “like” them. Nevertheless, such over-represenation of certain molecules, either originating from human cognitive biases, heuristics, social influences, from rationally made choices, from multiple isomeric instances, or a combination of all of the above, can lead to bias in databases.
Data de-duplication is the first step towards removing such bias, which reduces the concentration of the high density regions of the data manifold. In this case, de-duplication removes isomeric population information as well as repeated molecules across databases. The de-duplicated data is closer to a data manifold that has more homogenized density all over. Training on such a data manifold results in higher novelty in generations, as found for GP-MoLFormer-Uniq when compared to GP-MoLFormer, as shown in Table 6.
Although many existing molecular generative models trained on a much smaller and much focused datasets have demonstrated near-perfect (100%) novelty in generations, they are for most part not suitable for studying the trade-off between training data memorization and generation novelty. Investigating such phenomenon requires studying a generative chemical (language) model that has been trained on a broader-purpose and much larger dataset at scale. Our experiments attempt to address this under-explored aspect in this study. As shown in Table 6, novelty in GP-MoLFormer generations are lower compared to $\approx 100\%$ reported by baselines[14], but still sufficiently high for practical use. When compared with recent baselines, GP-MoLFormer generations are more dissimilar to test molecules (see the earlier sections and Table 1), though GP-MoLFormer’s test set is more diverse. And, finally, the low novelty in GP-MoLFormer’s generations is reflective of modeling its vast training set that represents the relative usage of molecules in real-world.
Similarly, the present study highlights the importance of studying generated sets of different sizes to obtain a comprehensive view of the quality of generations, particularly when the generative model is trained on data at scale. As GP-MoLFormer-Uniq aims to capture a training data manifold of more uniform density, which is enabled by de-duplicating the training SMILES, we see a 1% rise in novelty as we increase the number of generated samples from 30k to 10M. A similar observation has been reported in image generation [38] and language generation [39, 7]. To summarize, novelty in generations is influenced by the support provided by both the training distribution and the generated distribution, and should therefore be assessed relative to the sizes and diversity of those two sets.
These results in Table 6 complement and support earlier efforts focusing on studying scaling behaviors of chemical language models. One such noteworthy effort along this line is Frey et al. (2023)[40], where neural-scaling behavior in large chemical models was investigated by studying models with over 1B parameters and a scaling relation following a power law was established between training loss and model parameters. However, the models tested in that work were only pre-trained on datasets of size up to $\approx$ 10M data points, which is very small compared to the size of the chemical universe. In Ross et al. (2022)[13], the scaling behavior of MoLFormer, which is a transformer-based molecular encoder built using a masked language modeling objective, was studied. That work clearly established the scaling behavior underlying adaptation of a pre-trained model across downstream tasks, in which the number of model parameters was up to 47M while the number of training points considered was >1B. It was shown that a MoLFormer trained on 100M SMILES consistently underperformed across a wide variety of property prediction tasks, including quantum mechanical and physiological, when compared to the model trained on >1.1B SMILES, indicating predictive ability may benefit from such bias in training data. In contrast, the results in Table 6 show that a generative chemical language model trained on cleaner de-duplicated training data produces more novel generations.
We next investigate how these metrics change with varying number of generated and test molecules. Table 7 shows that, with increasing the generated pool size, scaffold similarity with respect to the test molecules becomes $>0.9$ while SNN reaches $>0.5$ when compared against 175k held-out test samples. When a larger test set of 1M molecules is used, further increases in both scaffold similarity and SNN are observed. These results imply that, with increasing size and diversity of the training data, the typical metrics used in assessing molecular generative models, such as various similarity measures with respect to a test set, should be carefully analyzed with generation and test sets that are larger in size compared to what is typically used in the field. Note that, even for 1M generations, GP-MoLFormer produces highly diverse molecules.
MOSES MoLFormer
[…]
175k 30k 0.9998 0.7383 0.5045 0.8655 0.0591 6.970 13.10 0.1844
100k 0.9998 0.8653 0.5045 0.8657 0.0279 6.967 13.10 0.1025
1M 0.9998 0.9375 0.5040 0.8658 0.0178 6.970 13.11 0.0741
1M 30k 0.9998 0.7702 0.5738 0.8655 0.0646 6.180 13.10 0.1684
100k 0.9998 0.9026 0.5740 0.8657 0.0331 6.179 13.10 0.0874
1M 0.9998 0.9786 0.5739 0.8658 0.0227 6.183 13.11 0.0600
Table 7: Investigation on the interplay between generation set size and test set size on modeling the molecular distribution of GP-MoLFormer-Uniq generations. Columns are the same as Table 1.
Scaling law for inference compute
We further attempt to fit a scaling law to the empirical trend observed in Table 6 for novelty with respect to generation size, which reflects the scale of inference compute. For both models, this trend appears to follow an exponential decay of the form:
$y=ae^{-bx}$ (1)
where $y$ is the novelty, $x$ is the generation size, and $a$ and $b$ are fitted parameters. In practice, $b$ is very close to 0 so we can rewrite this as:
$y=ae^{-10^{c}x}$ (2)
for ease of reading. These results can be further seen in Supplementary Figure S7. We observe that the fitted initial value $a$ is higher and decay constant $b$ is lower for GP-MoLFormer, meaning it starts with higher novelty and declines more slowly, when compared to GP-MoLFormer. To our knowledge, this is the first ever investigation of inference scaling of chemical language models.
@@ Conclusion
In this work, we investigate the effect of training data bias and diversity on the downstream performance of a generative chemical language model named GP-MoLFormer, built on top of a recent transformer architecture for chemical language modeling. We show the generality of the proposed GP-MoLFormer architecture on de novo generation and on two distinct targeted design tasks, i.e., scaffold-constrained molecular decoration and unconstrained property-guided molecular optimization, where the model performs better or on par when compared to the baselines. We further show how bias in training data can induce memorization, and thus impact the novelty of generations. We analyze how the commonly used metrics for comparing models’ generations with a held-out test set are affected by the diversity of the training distribution; hence, the sizes of the test set and the generation set should be carefully considered before making a conclusion based on those metrics. To our knowledge, this is the first report demonstrating training data memorization and its impact on downstream performance of a generative chemical language model pre-trained on billion-scale data. We also investigate effect of inference compute scaling and establish a scaling law between number of generations and novelty in them.
@@ Methods
The GP-MoLFormer decoder uses the transformer block used in MoLFormer. To avoid the quadratic complexity associated with regular attention computations in vanilla transformers [41], MoLFormer utilized a base 12-layer transformer architecture with linear attention [42], wherein each layer has $12$ attention heads and a hidden state size of $768$. A Generalized Random Feature map [43] for the linear attention was chosen.
@@
where $Q,K,V$ are the query, key, and value respectively, and $\varphi$ a random feature map. The GP-MoLFormer is trained on the next token prediction task using a cross-entropy objective: $L_{LM}(w1,...,wn)=\sum_{i}{\log{P}(w_{i}|w_{j<i})}$. Given the size of the transformer model and the efficient linear attention, GP-MoLFormer takes only around 3 milliseconds for a single forward pass during generation, using a single A100 GPU.
@@
We used two datasets for pre-training by combining SMILES from the PubChem [45] and the ZINC [9] databases with varying proportions from each. The dataset used in GP-MoLFormer training contains a total of 1.1B SMILES strings;  [^1] M of them are from the PubChem dataset, whereas the larger 1B portion comes from the ZINC database. The Uniq dataset is a de-duplicated version of that 1.1B size dataset, which comprises 650M SMILES. We utilized the tokenizer from Schwaller et al. (2019)[46] to construct a vocabulary. All SMILES sequences from both PubChem and ZINC are converted to a canonical format with no isomeric information using RDKit [47] followed by de-duplication (Uniq only) and tokenization. All unique tokens extracted from the resulting output give us a vocabulary of $2357$ tokens plus $5$ special tokens, resulting in a total of $2362$ vocabulary tokens which are used for all pre-trained models considered in this paper, irrespective of pre-training dataset size. The post tokenization sequence length of the molecules range from $1$ to just over $2000$ tokens. We decide to restrict the sequence length range from $1$ token to $202$ tokens, special tokens inclusive, to reduce computation time. Since over $99.4$ percent of all molecules from our dataset contain less than $202$ tokens, we hypothesize that the removal of molecules with more than $202$ tokens would be of minimal negative impact on pre-training.
Large-scale training and parallelization
For pre-training, we use the causal language model objective defined in Devlin et al. (2019)[48]. The training was performed for 4 epochs (for both 1.1B and 650M training dataset sizes) with a fixed learning rate of $1.6\times 10^{-4}$ and a batch size of $1600$ molecules per GPU on a total of $16$ A100 80 GB GPUs over $2$ servers connected via EDR Infiniband fabric. It should be noted that as the number of GPUs utilized increased, we found an increase in learning rate was necessary up to a factor of $8$. The GP-MoLFormer model was trained for 28.75 hours per epoch, for 115 hours total training time, while the GP-MoLFormer-Uniq model was trained for 19.75 hours per epoch, less than 80 hours total training time.
@@
For evaluation, we do a forward pass using the following sequence: the first 20 prompt embeddings + the <bos> embedding + the input molecule embeddings + the <unk> embedding. After that, we sample from the token distribution generated by GP-MoLFormer until <eos> is encountered. For all pair-tuning experiments, batch size was set to 35, the learning rate was fixed at $3\times 10^{-2}$, and the number of epochs run was 1000. Each epoch took 6 minutes to complete on a single GPU.
Author contributions
J.R. and B.B. trained the GP-MoLFormer models. S.H. and J.R. performed pair-tuning experiments. S.H., V.C, J.R., and J.N. conducted the model evaluation experiments. P.D., Y.M., J.R., B.B., V.C, and S.H. contributed to the idea generation, study design, result analyses, and paper writing.
Conflicts of interest
There are no conflicts to declare.
Acknowledgments
Authors acknowledge IBM Research for their support.